adorenaysha:

A sequence is defined recursively by the formula f(n + 1) = f(n) + 3 . The first term of the sequence is –4. What is the next term in the sequence?

dude:

Substitute the first value into the equation \(f(n + 1) = f(n) + 3 \rightarrow f(n + 1) = -4 + 3\)